# An empirical appraisal of methods for the dynamic prediction of survival with numerous longitudinal predictors

### v1

@@ 4.3 PBC2
Table 3: Cross-validated C index estimates (and standard deviation) for the PBC2 dataset.
@@
Time-dependent AUC. PRC achieves the highest tdAUC values at all landmarks. At landmark 2 it is followed by the Static Cox model, DynForest and LOCF landmarking; FunRSF and MFPCCox have considerably lower predictive accuracy. At landmark 3, DynForest, LOCF landmarking and MFPCCox achieve similar tdAUC values. At landmark 3.5, the trends are somewhat harder to extrapolate. It’s interesting to note that at all landmarks, the performance of the static Cox model deteriorates quickly as the prediction horizon increases. Moreover, for the first time we can observe that at landmark 2.5, FunRSF outperforms MFPCCox.
@@
Brier score. At landmark 2.5, the estimated Brier scores for landmarking, static Cox, PRC and DynForest are very similar, specially at the earlier prediction horizons. The two methods that rely on MFPCA are once again outperformed by all other methods. At landmark 3, the static Cox model has lower Brier score for predictions after 1 and 2 years from the landmark, and PRC for predictions from 3 years from the landmark onwards. At landmark 3.5, the differences between methods tend to be small for horizons at 4.5 and 5.5 years, after which the lowest Brier scores are achieved by landmarking and PRC, and the highest by FunRSF and the static Cox model.
Summary.The estimates of predictive performance for the PBC2 study are generally less accurate than for the ADNI and ROSMAP datasets due to the smaller sample size. Despite this, the results are somewhat similar to those obtained on the previous two studies.
@@ 4.4 Computing time
To give an idea of how much computing time differs across methods, the average computing time required to estimate each model in $k-1$ folds and to obtain predictions for the remaining fold during the RCV process is shown in Supplementary Tables 11-13. It can be observed that computing time typically increases with the sample size, and slightly decreases as the landmark increases (because less individuals are still at risk at later landmark times).
Estimation of models that do not model the longitudinal data, i.e. the static Cox and LOCF landmarking models, is almost immediate: this is due to the fact that these approaches only require to estimate a Cox model through R’s function coxph( ). Among the other methods, MFPCCox is the fastest method, followed by FunRSF: both methods can be estimated in less than a minute on all datasets. PRC and DynForest are more computationally intensive: on average, estimating PRC takes 0.57 minutes on ADNI, 1.65 minutes on ROSMAP and 7.7 seconds on PBC2; DynForest takes almost 10 minutes on ADNI, 21.1 minutes on ROSMAP and 2.5 minutes on PBC2.
@@ 5 Discussion
In this study we reviewed and evaluated the predictive performance of four recently-proposed dynamic prediction methods (MFPCCox, PRC, FunRSF, and DynForest) that are capable of using numerous longitudinal predictors to predict survival. Each method uses a different combination of MFPCA or LMMs to model the longitudinal predictors, and of Cox model or RSF to predict the survival outcome: therefore, by comparing these methods we can also get a feeling of the extent to which the choices of using MFPCA or LMMs, and a Cox model or a RSF, can affect predictive performance. These 4 methods were further compared to a static Cox model that does not use any information after baseline, and to the LOCF landmarking approach.
Our results show that methods that use LMMs (PRC and DynForest) clearly outperform methods that use MFPCA (MFPCCox and FunRSF) to model the trajectories of the longitudinal covariates. PRC is frequently the best performing method with respect to the different metrics considered,; it is usually followed by DynForest and MFPCCox, while FunRSF almost always exhibits the worst predictive performance.
As concerns the method used to predict survival, it seems that on the datasets considered in this study, using a Cox model allows to achieve a better predictive performance than using a RSF, as highlighted by the fact that conditionally on the same method used to model the longitudinal predictors, PRC outperforms DynForest and MFPCCox outperforms FunRSF. Interestingly, LOCF landmarking often showed a good predictive performance, pointing out how such a simple and intuitive prediction method can often deliver good predictions of survival.
When considering computing time, most methods proved to be quite fast to estimate. DynForest was by far the most time-consuming method: this can be attributed to the fact that instead of estimating LMMs once, DynForest re-estimates the LMMs after each node split. Besides increasing computing time, this can also lead to more frequent convergence problems. Another method that took longer was PRC, showing how in general, estimation of LMMs is more time-consuming than estimation of the MFPCA scores. Nevertheless, using LMMs clearly resulted in higher predictive accuracy, justifying the increased computing time.
The main advantage of the two-step methods considered in this study is that they make it possible to extend the applicability of dynamic prediction to studies with tens, hundreds or (potentially) even hundreds of longitudinal covariates. At the same time, these new methods lack some elements of flexibility:
  • To be able to estimate the MFPCA scores, with MFPCCox and FunRSF it was necessary to standardize the measurement times across subjects so as to have a shared time grid. This is undesirable, both for longitudinal studies with fixed measurement times, where individual visits may deviate from the planned date, and for longitudinal studies where there is no planned visit schedule;
  • DynForest and PRC are limited to LMMs. This induced us to apply skewness-reducing transformations to strongly-skewed longitudinal predictors so as to reduce convergence problems; moreover, in the case of discrete predictors it would be preferable to resort to GLMMs;
@@
  • none of the 4 methods currently allows to deal with interval censoring, which is common in longitudinal studies;
  • MFPCCox, FunRSF and PRC don’t allow to model competing risks.
All four methods require users to make choices about tuning parameters. This is less of a problem with PRC, that uses cross-validation to select the value(s) of the penalty parameter(s), and with MFPCCox, where one only needs to choose the percentage of variance explained that the MFPCA scores need to retain. In addition to the percentage of variance explained, FunRSF also requires the specification of the number of variables to choose as candidates for node splitting, and of the minimal number of subjects that a terminal node needs to contain. DynForest requires users to specify three tuning parameters: number of candidate covariates, and number of subjects and of events that determine the termination of node splitting.
@@
Finally, PRC and DynForest come with software implementations (Signorelli,, 2023; Devaux et al.,, 2023) that make it easy for users to implement those methods. MFPCCox and FunRSF, instead, lack a proper software implementation, forcing users to adapt the code presented in the corresponding publications to their own datasets.
Supplementary tables and figures
@@
Supplementary tables and figures are available in the “Supplementary Material” pdf file.
Data availability
ROSMAP. The data from the ROSMAP study are available upon motivated request from the Rush Alzheimer’s Disease Center (RADC) of the RUSH University. ROSMAP resources can be requested at https://www.radc.rush.edu and www.synpase.org. All ROSMAP participants enrolled without known dementia and agreed to detailed clinical evaluation and brain donation at death (Bennett et al.,, 2018). The ROS and MAP studies were approved by an Institutional Review Board of Rush University Medical Center. Each participant signed an informed consent, Anatomic Gift Act, and repository consent to allow their data to be repurposed.
ADNI. The ADNI data are available upon motivated request from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (http://adni.loni.usc.edu).
PBC2. The PBC2 dataset is publicly available and can be found in several R packages including, among others, DynForest, joineRML and pencal.

### v2

@@ 4.3 Predictive performance on the PBC2 data
Table 3: Cross-validated C index estimates (and standard deviation) for the PBC2 dataset. For each landmark, the top two methods are highlighted in bold.
C index (Table 3). At all landmark times, PRC achieves the highest value of the C index. Interestingly, the static Cox model is the second best performing method at landmarks 2.5 and 3 and third best at landmark 3.5, suggesting that the added predictive value of the longitudinal measurements may be more limited for the PBC2 dataset. LOCF landmarking is the third bist method at landmarks 2.5 and 3, and second best at landmark 3.5. It is followed by DynForest, MFPCCox and FunRSF. By comparing the standard errors in Tables 1, 2 and 3, we can see that the standard errors around the estimated C index are larger on PBC2 than on ADNI and ROSMAP. This result can be ascribed to the substantially smaller sample size of this dataset.
@@
Time-dependent AUC (Figure 5). PRC achieves the highest tdAUC values at all landmarks. At landmark 2, it is followed by the Static Cox model, DynForest and LOCF landmarking; FunRSF and MFPCCox have considerably lower predictive accuracy. At landmark 3, DynForest, LOCF landmarking and MFPCCox achieve similar tdAUC values. At landmark 3.5, the trends are somewhat harder to extrapolate. Lastly, at all landmarks the performance of the static Cox model deteriorates quickly as the prediction horizon increases.
@@
Brier score (Figure 6). At landmark 2.5, the estimated Brier scores for landmarking, static Cox, PRC and DynForest are very similar, especially at the earlier prediction horizons. The two methods that rely on MFPCA are once again outperformed by all other methods. At landmark 3, the static Cox model has lower Brier score for predictions after 1 and 2 years from the landmark, and PRC for predictions from 3 years from the landmark onwards. At landmark 3.5, the differences between methods tend to be small for horizons at 4.5 and 5.5 years, after which the lowest Brier scores are achieved by landmarking and PRC, and the highest by FunRSF and the static Cox model.
Summary.The estimates of predictive performance for the PBC2 study are generally less accurate than for the ADNI and ROSMAP datasets due to the smaller sample size. Despite this, the relative performance of the different methods is to a wide extent in line with the patterns observed on the ADNI and ROSMAP datasets.
@@ 4.4 Computing time
We now turn our attention to computing time. Tables 4, 5, and 6 report the average computing time required to estimate each model in $k-1$ folds and to obtain predictions for the remaining fold (averaged over all RCV iterations). Computing time is generally higher for the ROSMAP dataset, where we have the largest sample size and number of predictors, and lowest for the PBC2 dataset where $n$, $P$ and $Q$ are small. By comparing computing time across landmarks, we observe that the computing time typically decreases as the landmark increases; this is simply an artifact of the fact that at later landmarks less individuals are still at risk, reducing the sample size used for model training and prediction.
Estimation of models that do not model the longitudinal data, i.e. the static Cox and LOCF landmarking models, is almost immediate: this is because these approaches only require to estimate a Cox model through R’s function coxph( ). The multi-step methods are more computationally intensive. Among them, MFPCCox is the fastest method, followed by FunRSF. This shows that multi-step methods that use MFPCA are faster than those that use LMMs. The faster computing time of MFPCCox is likely due to the use of a Cox model, whose estimation is faster than that of RSF (used by FunRSF). Lastly, we can observe that among the LMM-based methods, PRC is substantially faster (between 12 and 20 times, depending on the dataset) than DynForest. This last result is mainly due to the fact that in DynForest, the LMMs are re-estimated at each node split within each random survival tree, making this method considerably slower than all other alternatives.
Landmark
Method 2 3 4 Average
Static Cox 0.009 0.007 0.006 0.007
Landmarking 0.010 0.007 0.006 0.008
MFPCCox 0.080 0.046 0.046 0.057
PRC 0.776 0.482 0.453 0.571
FunRSF 0.240 0.122 0.125 0.163
DynForest 12.501 9.077 8.099 9.892
Table 4: Average computing time per CV fold (in minutes) for the ADNI dataset.
Landmark
Method 2 3 4 5 6 Average
Static Cox 0.022 0.019 0.016 0.014 0.011 0.017
Landmarking 0.023 0.020 0.017 0.014 0.011 0.017
MFPCCox 0.194 0.135 0.066 - 0.064 0.115
PRC 3.755 1.151 1.138 1.115 1.124 1.657
FunRSF 0.604 0.367 0.115 - 0.124 0.302
DynForest - 11.504 17.657 - 34.147 21.102
Table 5: Average computing time per CV fold (in minutes) for the ROSMAP dataset.
Landmark
Method 2.5 3 3.5 Average
Static Cox 0.253 0.246 0.177 0.226
Landmarking 0.260 0.240 0.188 0.229
MFPCCox 0.902 0.522 0.556 0.660
PRC 7.832 7.616 7.762 7.737
FunRSF 2.776 1.393 1.526 1.898
DynForest 186.136 136.747 128.914 150.599
Table 6: Average computing time per CV fold (in seconds) for the PBC2 dataset.
@@ 5 Discussion
This study aimed to benchmark four recently-proposed multi-step dynamic prediction methods (MFPCCox, PRC, FunRSF, and DynForest) that are capable of using numerous longitudinal predictors to predict survival. Each method uses a different combination of MFPCA or LMMs to model the longitudinal predictors, and of Cox models or RSF to predict the survival outcome: therefore, by comparing these methods we can also get a feeling of the extent to which the choices of using MFPCA or LMMs, and a Cox model or an RSF, can affect predictive performance. Hereafter we summarize the findings and limitations of this study, focusing our attention on the modelling flexibility, predictive performance, and ease of use of the different methods.
5.1 Modelling flexibility
The four multi-step methods reviewed in this article represent an important and valuable addition to the dynamic prediction toolbox. Differently from traditional landmarking methods, multi-step dynamic prediction methods employ models that enable efficient modelling of the longitudinal data, and to account for the presence of measurement errors. These methods are particularly suitable for datasets where where tens, hundreds or (potentially) thousands of longitudinal predictors may be available, where the estimation of JMs is not only very computationally intensive, but usually also unfeasible.
At the same time, when modelling real-world longitudinal and survival data it may sometimes necessary to deal with competing risks, interval censoring, and missing data. Each of these features can lead to complications with some of the multi-step methods considered in this work, at least in their present state of development:
  1. 1.
    whereas DynForest has been developed within a competing risks framework, MFPCCox, FunRSF and PRC don’t allow to model competing risks directly. While extending them to competing risks is possible, currently users would need to do that by themselves;
  2. 2.
@@ 5.1 Modelling flexibility
    none of the four methods currently allows interval censored survival outcomes, which are common in longitudinal studies. Therefore, users need to simplify the survival outcome to right-censored;
  3. 3.
    PRC and DynForest are currently limited to LMMs (and multivariate LMMs in the case of PRC). Extending them to GLMMs would enable a more appropriate modelling of skewed as well as discrete longitudinal covariates, although it can be anticipated that such an extension would be computationally more complex, potentially leading to more frequent problems during the estimation of the GLMMs and the computation of the predicted random effects;
  4. 4.
    as seen on the ROSMAP dataset, a high percentage of missingness in the longitudinal data can considerably complicate the estimation of MFPCCox, FunRSF and DynForest. For MFPCCox and FunRSF, this makes it harder to estimate the covariance matrices during the univariate FPCA steps. For DynForest, this creates problems with the estimation of LMMs in children nodes with a small number of subjects.
5.2 Predictive performance
The results presented in Section 4 are mostly consistent across datasets and performance measures. The most interesting patterns are:
  1. 1.
    multi-step methods that use LMMs (PRC and DynForest) clearly outperform methods that use MFPCA (MFPCCox and FunRSF) to model the trajectories of the longitudinal covariates. PRC is frequently the best performing method with respect to the different metrics considered; it is usually followed by DynForest and MFPCCox, while FunRSF almost always exhibits the worst predictive performance;
  2. 2.
    conditionally on the same method being used to summarize the longitudinal covariates, the methods that use a Cox model (PRC and MFPCCox) tend to outperform the corresponding method that uses RSF (DynForest and FunRSF, respectively). However, the difference is substantially smaller than the one between LMM-based and MFPCA-based methods, and it may be dataset specific, as the performance of RSF may improve on datasets with larger sample sizes;
  3. 3.
    interestingly, LOCF landmarking often showed a good predictive performance, often outperforming some of the multi-step methods. This suggests that this simple and intuitive prediction method can often deliver good predictions of survival. Thus, we recommend that practitioners interested in the use of multi-step methodsir always compare the predictive performance to that of LOCF landmarking, because the latter approach may be preferable in case of similar predictive performance;
  4. 4.
    the extent to which longitudinal data may improve the predictive accuracy of a dynamic RPM over a static RPM varies across datasets. To quantify the extent of this improvement for a specific dataset, we suggest comparing the predictive performance of the developed dynamic RPM to that of a simpler static RPM that only uses baseline information.
5.3 Software and computing time
Software availability and computing time are two matters of practical importance when attempting to estimate and implement a RPM, affecting the ease with which a practicioner may be able to estimate a RPM and employ it to compute predictions.
@@ 5.3 Software and computing time
PRC and DynForest come with fairly advanced software implementations and documentation (Signorelli, , 2024; Devaux et al., , 2024) that make it easier for users to implement those methods. On the contrary, MFPCCox and FunRSF lack software implementations and documentation: this forces users to reuse the code presented in the corresponding publications, and often requires them to adapt it to deal with missing data and irregular measurement grids.
MFPCCox and FunRSF are computationally inexpensive: their estimation takes less than a minute on all three datasets considered in our benchmarking. PRC is a bit more intensive, taking less than a minute on PBC2 and ADNI, and approximately 1.66 minutes on ROSMAP. Finally, DynForest is by far the slowest method: its estimation takes about 2.5 minutes on PBC2, 10 on ADNI, and 21 on ROSMAP.
Supplementary material and R code
@@
Supplementary tables and figures are available in the “Supplementary Material” pdf file. The R code used to perform the analyses described in this paper is available at
https://github.com/mirkosignorelli/comparisonDynamicPred.